# Cooperative Bearing-Only Target Pursuit via Multiagent Reinforcement Learning: Design and Experiment

### v1

@@ I Introduction
The pursuit-evasion game has been a subject of significant research focus over several decades. It serves as a fundamental yet versatile mathematical framework with diverse practical applications, including autonomous guidance [1], criminal pursuit [2], aerospace [3] and robotics [4]. In a multiagent pursuit-evasion game, an extension of the classic one-on-one pursuit-evasion scenario, multiple pursuers coordinate to capture or intercept evaders within a given environment (see Fig. 1). Understanding and analyzing multiagent pursuit-evasion games provide valuable insights into cooperative strategic decision-making. However, solving multiagent pursuit-evasion games is a complex challenge due to unknown target states, dynamic interactions and multiple constraints among agents, resulting in a high-dimensional, non-linear, and often non-convex problem.
@@
The first step to realize unknown target pursuit is to estimate target state. Vision-based solutions serve as a prominent method for target perception. However, the readily available information from visual sensors is target bearing, and depth can not be recovered directly when the target size is unknown [5, 6]. Although SLAM-based methods offer a potential solution, they rely on feature points extraction or 3D bounding boxes for state estimation [7] [8]. These approaches become ineffective when the target is distant and small, as insufficient stable features can be extracted, and 3D detection is not feasible [9]. These limitations restrict us to addressing the bearing-only target state estimation problem [10, 9]. A notable challenge in bearing-only state estimation lies in the inherent non-linearity between the measurement and target state, which leads to biased or even divergent estimation outcomes for conventional methods like the extended Kalman filter. Although the utilization of modified polar or spherical coordinates[11, 12], and the pseudo-linear Kalman filter (PLKF) [13, 14, 15] alleviates the divergence problem, most of them are analyzed in 2D scenarios. When applied to 3D cases, the two-angle representation including azimuth and elevation introduces singularities at an elevation of 90°$\degree$.
@@ III Target State Estimation
Effective target pursuit requires first estimating the target state. In this section, we introduce a novel 3D bearing-only cooperative pseudo-linear information filter (PLIF) to estimate the target position and velocity from multiple noisy bearing measurements.
@@ III-B Measurement Equation
where $H_{\_}{i,k}=\begin{bmatrix}P_{\_}{\tilde{\lambda}_{\_}{i,k}}&0_{\_}{3\times 3}\end{bmatrix}$, $V_{\_}{i,k}=\hat{r}_{\_}{i,k}^{-}P_{\_}{\tilde{\lambda}_{\_}{i,k}}$; $\hat{r}_{\_}{i,k}^{-}$ is simply the norm of the first three components of $\hat{x}_{\_}{i,k}^{-}$; $\hat{y}_{\_}k$ and $Y_{\_}k$ are the posterior information state estimate and information matrix. The reason for using pseudo-inverse $\dagger$ is that the matrix $(V_{\_}{i,k}\Sigma_{\_}iV_{\_}{i,k}^{T})$ is rank deficient since $P_{\_}{\tilde{\lambda}_{\_}{i,k}}$ in $H_{\_}{i,k}$ is rank deficient. The is a common practice when the inverse does not exist [27, 28, 29]. Following the naming convention in [30, 10], The proposed cooperative uniform pseudo-linear information filter is coined as u-PLIF.
@@ IV-A Environment, Agent Dynamics and Action Space
We first developed a physics-based simulation environment for MARL, where pursuer and target vehicles interact within a two-dimensional continuous space, as shown in Fig. 3a. Pursuer agents and their FoV are in red, green and blue, respectively. The target is depicted in orange with a circle representing the control range. Additional details of an agent are illustrated in Fig. 3b. A body-frame $\{b\}$ is rigidly attached on the agent with its x-axis $x_{\_}b$ aligned with the forward direction of the vehicle. A simulated forward-looking camera with limited FoV is assumed to be installed on the agent aligned with the x-axis of the body frame. An agent is subject to three actions: the angular velocity of the heading $v_{\_}h\in\mathbb{R}$, and the linear velocity commands $v_{\_}x,v_{\_}y\in\mathbb{R}$ in the x-, y-axis of $\{b\}$. To summarize, the actions of the agent are $[v_{\_}h,v_{\_}x,v_{\_}y]^{T}\in\mathbb{R}^{3}$.
@@
where Eq. (6a) and (6b) describe the translational dynamics, $p\in\mathbb{R}^{2}$ is the position, $v\in\mathbb{R}^{2}$ is the velocity, $a\in\mathbb{R}^{2}$ is the linear acceleration control input, $f_{\_}a\in\mathbb{R}^{2}$ is the elastic force between contact agents, $f_{\_}b\in\mathbb{R}^{2}$ is the elastic force between agents and boundaries. Elastic forces reflect physical collision dynamics following Hooke’s law, and sum up when an agent contacts multiple other agents or boundaries as $f_{\_}a=\sum_{\_}jf_{\_}{a,j}$ and $f_{\_}b=\sum_{\_}jf_{\_}{b,j}$, and $m\in\mathbb{R}^{+}$ is the mass. Equations (6c) and (6d) describe the rotational dynamics, where $R\in\mathbb{R}^{2\times 2}$ is the rotation matrix from the vehicle’s body frame $\{b\}$ to the inertial frame $\{i\}$, $\omega\in\mathbb{R}$ is the angular speed. Specifically,
@@ IV-C Sim-to-Real
Figure 4: Experiment Result (See the accompanying video at https://youtu.be/HO7FJyZiJ3E.)

### v2

@@ I Introduction
The pursuit-evasion game has been a subject of significant research focus over several decades. It serves as a fundamental yet versatile mathematical framework with diverse practical applications, including autonomous guidance [1], criminal pursuit [2], aerospace [3] and robotics [4]. In a multiagent pursuit-evasion game, an extension of the classic one-on-one pursuit-evasion scenario, multiple pursuers coordinate to capture or intercept evaders within a given environment (see Fig. 1). Understanding and analyzing multiagent pursuit-evasion games provide valuable insights into cooperative strategic decision-making [5]. However, solving multiagent pursuit-evasion games is a complex challenge due to unknown target states, dynamic interactions and multiple constraints among agents, resulting in a high-dimensional, non-linear, and often non-convex problem.
@@
The first step to realize unknown target pursuit is to estimate target state. Vision-based solutions serve as a prominent method for target perception. However, the readily available information from visual sensors is target bearing, and depth can not be recovered directly when the target size is unknown [6, 7, 8]. Although SLAM-based methods offer a potential solution, they rely on feature points extraction or 3D bounding boxes for state estimation [9] [10]. These approaches become ineffective when the target is distant and small, as insufficient stable features can be extracted, and 3D detection is not feasible [11]. These limitations restrict us to addressing the bearing-only target state estimation problem [12, 11]. A notable challenge in bearing-only state estimation lies in the inherent non-linearity between the measurement and target state, which leads to biased or even divergent estimation outcomes for conventional methods like the extended Kalman filter. Although the utilization of modified polar or spherical coordinates[13, 14], and the pseudo-linear Kalman filter (PLKF) [15, 16, 17] alleviates the divergence problem, most of them are analyzed in 2D scenarios. When applied to 3D cases, the two-angle representation including azimuth and elevation introduces singularities at an elevation of 90°$\degree$.
@@ III Target State Estimation
Effective target pursuit requires first estimating the target state. In this section, we extend the work in [12], and introduce a novel 3D bearing-only cooperative pseudo-linear information filter (PLIF) to estimate the target position and velocity from multiple noisy bearing measurements.
@@ III-B Measurement Equation
where $H_{\_}{i,k}=\begin{bmatrix}P_{\_}{\tilde{\lambda}_{\_}{i,k}}&0_{\_}{3\times 3}\end{bmatrix}$, $V_{\_}{i,k}=\hat{r}_{\_}{i,k}^{-}P_{\_}{\tilde{\lambda}_{\_}{i,k}}$; $\hat{r}_{\_}{i,k}^{-}$ is simply the norm of the first three components of $\hat{x}_{\_}{i,k}^{-}$; $\hat{y}_{\_}k$ and $Y_{\_}k$ are the posterior information state estimate and information matrix. The reason for using pseudo-inverse $\dagger$ is that the matrix $(V_{\_}{i,k}\Sigma_{\_}iV_{\_}{i,k}^{T})$ is rank deficient since $P_{\_}{\tilde{\lambda}_{\_}{i,k}}$ in $H_{\_}{i,k}$ is rank deficient. This is a common practice when the inverse does not exist [29, 30, 31]. Following the naming convention in [32, 12], The proposed cooperative uniform pseudo-linear information filter is coined as u-PLIF.
@@ IV-A Environment, Agent Dynamics and Action Space
We first developed a physics-based simulation environment for MARL, where pursuer and target vehicles interact within a two-dimensional continuous space [5], as shown in Fig. 3a. Pursuer agents and their FoV are in red, green and blue, respectively. The target is depicted in orange with a circle representing the control range. Additional details of an agent are illustrated in Fig. 3b. A body-frame $\{b\}$ is rigidly attached on the agent with its x-axis $x_{\_}b$ aligned with the forward direction of the vehicle. A simulated forward-looking camera with limited FoV is assumed to be installed on the agent aligned with the x-axis of the body frame. An agent is subject to three actions: the angular velocity of the heading $v_{\_}h\in\mathbb{R}$, and the linear velocity commands $v_{\_}x,v_{\_}y\in\mathbb{R}$ in the x-, y-axis of $\{b\}$. To summarize, the actions of the agent are $[v_{\_}h,v_{\_}x,v_{\_}y]^{T}\in\mathbb{R}^{3}$.
@@
where Eq. (6a) and (6b) describe the translational dynamics, $p\in\mathbb{R}^{2}$ is the position, $v\in\mathbb{R}^{2}$ is the velocity, $a\in\mathbb{R}^{2}$ is the linear acceleration control input, $f_{\_}a\in\mathbb{R}^{2}$ is the elastic force between contact agents, $f_{\_}b\in\mathbb{R}^{2}$ is the elastic force between agents and boundaries. Elastic forces reflect physical collision dynamics following Hooke’s law, and sum up when an agent contacts multiple other agents or boundaries as $f_{\_}a=\sum_{\_}jf_{\_}{a,j}$ and $f_{\_}b=\sum_{\_}jf_{\_}{b,j}$, and $m\in\mathbb{R}^{+}$ is the mass. Equations (6c) and (6d) describe the rotational dynamics, where $R\in\mathbb{R}^{2\times 2}$ is the rotation matrix from the vehicle’s body frame $\{b\}$ to the fixed inertial frame $\{i\}$, $\omega\in\mathbb{R}$ is the angular speed. Specifically,
@@ IV-C Sim-to-Real
Figure 4: Experiment Result (See the accompanying video at https://youtu.be/HO7FJyZiJ3E. More original experimental footage is available at https://youtu.be/70ead3rQcok.)